# Generative Modeling with Explicit Memory

### v1

@@ 4.1 External Memory Bank
Inspired by previous studies [27], which demonstrated that generalization in diffusion models arises from geometry-adaptive harmonic representations, we aim to design a memory bank that: (a) Provides sufficient semantic and representational information to assist the diffusion model in generating high-quality, realistic images that align with the true data distribution. (b) Avoids excessive detailed or textual information to prevent overfitting to the training data and ensure robust generalization.
We define the memory bank for diffusion models as a matrix $\mathbf{M}\in\mathbb{R}^{n\times m}$, encompassing $n$ memory snippets. Each snippet $\mathbf{s}$ contains $m$ units and satisfies $\|\mathbf{s}\|=1$:
$\mathbf{M}=\begin{bmatrix}\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{n}\end{bmatrix}^{\top}\,.$ (6)
We employ a robust representation model $\boldsymbol{f}$ under the condition that for every input $\mathbf{x}$ sampled from data distribution $D$, the produced normalized feature $\nicefrac{{\boldsymbol{f}(\mathbf{x})}}{{\|\boldsymbol{f}(\mathbf{x})\|}}$ always corresponds to a vector $\mathbf{s}$ that exists in the set $\mathbf{M}$. In conclusion, our goal is to ensure that all semantic information contained in $D$ is captured by $\mathbf{M}$.
To achieve this goal, we formalize the optimization objective for $\mathbf{M}$ as follows (see our analysis in Appendix A.1):
$\operatorname*{arg\,min\,}_{\mathbf{M}\in\mathbb{R}^{n\times m}}\left\{\left(\mathbb{E}_{\mathbf{x}\sim D}\left[\max\left\{\mathbf{M}\frac{\boldsymbol{f}(\mathbf{x})}{\|\boldsymbol{f}(\mathbf{x})\|}\right\}-1\right]\right)^{2}\right\}\,,$ (7)
where $\max(\cdot)$ denotes the maximum value of a vector. Note that $n$ and $m$ are predefined hyperparameters balancing the memorization capacity of $\mathbf{M}$ against storage constraints. Specifically, increasing $n$ and $m$ enhances the memorization capacity of $\mathbf{M}$ but also increases storage requirements. Detailed settings and analyses of $n$ and $m$ are provided in Appendix A.2.
Usage of representation models for building the memory bank.
Our memory bank framework is adaptable to various representation models $\boldsymbol{f}$, each with distinct capabilities in capturing image information. This variability impacts the performance of the trained generative models. Therefore, in this study, we utilize self-supervised representation models based on the following reasoning:
@@
    Previous research [70] demonstrates that self-supervised models can expedite the training of diffusion models, which can be seamlessly integrated with a memory bank.
@@
    Prior work [3, 74, 62] points out that self-supervised models capture semantic information more effectively than supervised alternatives.
We rigorously assess the effectiveness of different representation models within our framework to build the memory bank like CLIP visual encoder [54] and BERT [11] (see Appendix A.3 and Appendix F).
@@ 4.2 Training Objective
Our proposed memory bank is designed to supply semantic information about the data distribution, supporting both the training and inference phases of diffusion models. Therefore, we incorporate semantic information into diffusion models during the training phase by slightly modifying the diffusion model loss function as follows:
@@
where $\mathbf{x}_{0}\sim D,\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),\dot{\alpha}_{t}=\frac{\mathrm{d}\alpha_{t}}{\mathrm{d}t},\dot{\sigma}_{t}=\frac{\mathrm{d}\sigma_{t}}{\mathrm{d}t},\lambda\sim p(\lambda)$ and $\mathbf{v}_{\boldsymbol{\theta}}$ is the velocity estimated by SiT backbone. This approach trains diffusion models using a interpolant way over various noise levels [44], incorporating an additional input snippet $\mathbf{s}$, which was selected from memory bank $\mathbf{M}$ using cosine similarity:
$\mathbf{s}=\operatorname*{arg\,max\,}_{\mathbf{s}_{i}\in\mathbf{M}}\left\{\frac{\boldsymbol{f}(\mathbf{x}_{0})}{\|\boldsymbol{f}(\mathbf{x}_{0})\|}\cdot\mathbf{s}_{i}\right\}\,.$ (9)
To ensure a fair comparison with the baselines, we use the exact same parameter settings of $\alpha_{t}$ and $\sigma_{t}$ as in previous studies [44, 70]. For a detailed description of the training and generation processes, please refer to the framework outlined in Algorithm 1.
Multiple usage of representation models for enhancing training diffusion models.
A very recent work REPA [70] introduces a regularization term that leverages representation models within the diffusion model’s loss function to enhance training efficiency. Therefore, we can integrate this regularization with our modified diffusion loss (8):
$\mathcal{L}(\boldsymbol{\theta})\leftarrow\mathcal{L}(\boldsymbol{\theta})+\mathcal{L}_{\textsc{REPA}}(\boldsymbol{\theta})$ (10)
$\text{with}\;\;\mathcal{L}_{\textsc{REPA}}(\boldsymbol{\theta}):=-\mathbb{E}_{\mathbf{x}_{0},t}$ $\left[\operatorname{sim}\left(\boldsymbol{f}(\mathbf{x}_{0}),\boldsymbol{\psi}(\mathbf{h})\right)\right]\,,$
where $\boldsymbol{f}(\mathbf{x}_{0})$ denotes the pretrained representation, and $\boldsymbol{\psi}(\mathbf{h})$ represents the projection of the intermediate output of diffusion transformer block $\mathbf{h}$ using a trainable projection head $\boldsymbol{\psi}$, implemented as a multilayer perceptron (MLP).
@@ 4.3 Sampling with Memory Bank
The sampling stage of unconditional generation in diffusion models [24, 42] aims to transform Gaussian noise into the true data distribution within an end-to-end framework. This iterative process employs SDE or ODE solvers (refer to Section 3) to progressively convert the noise into the desired data distribution. Our framework GMem with the memory bank also follows this end-to-end generation process, namely, feeding the noise into the neural network, together with the corresponding memory snippet $\mathbf{s}$.
Specifically, our GMem transforms the noise distribution into a corresponding bank snippet index distribution (a.k.a., noise-to-index transformation) and samples indices from this distribution to obtain memory snippets. We propose a two-step transformation methodology below—namely (1) normal distribution to uniform distribution and (2) uniform distribution to index distribution—that rigorously converts Gaussian-distributed noise into the index distribution. Such a strategy allows the index distribution exhibiting two key characteristics: firstly, similar noise inputs should correspond to similar indices, thereby maintaining coherence in the selection process; secondly, each generated index should possess an equal probability of being selected during the training phase, promoting unbiased learning.
Step 1: normal distribution to uniform distribution.
The initial step leverages the Probability Integral Transform [52] to convert Gaussian-distributed noise into a uniformly distributed variable. Specifically, we begin with a high-dimensional noise vector $\boldsymbol{\epsilon}\in\mathbb{R}^{d}$ that follows a Gaussian distribution. By computing and scaling the sum of the vector, we obtain a one-dimensional Gaussian variable $\overline{\epsilon}=\frac{1}{\sqrt{d}}\sum_{i=1}^{d}\epsilon_{i}$, which follows the standard normal distribution $\mathcal{N}(0,1)$. The cumulative distribution function of this Gaussian variable, denoted by $\Phi(\overline{\epsilon})$, transforms the Gaussian noise into a uniform distribution over the interval $[0,1]$ and can be mathematically expressed as:
[…]
Step 2: uniform distribution to index distribution.
Given that the memory bank size may be much smaller than the total number of images in the dataset, a direct uniform distribution may not accurately reflect the true index distribution. To address this, we first record the index distribution from the training set, denoted as $\mathbf{P}_{\text{index}}=[p_{1},\ldots,p_{i},\ldots,p_{n}]$, where each $p_{i}$ represents the probability of selecting the $i$-th snippet $\mathbf{s}_{i}$ during training. Utilizing inverse transform sampling [52] , we map the uniform variable to the index distribution by defining the variable $k$ as the smallest index $i$ for which the cumulative probability $\sum_{j=1}^{i}p_{j}$ exceeds the uniform variable:
$k=\min\left\{i\in\{1,\ldots,n\}\ \bigg{|}\ \Phi(\overline{\epsilon})\leq\sum_{j=1}^{i}p_{j}\right\}\,.$
Upon completion of aforementioned noise-to-index transformation, the original noise vector $\boldsymbol{\epsilon}$ and the selected memory snippet $\mathbf{s}_{k}$ are combined and fed into the diffusion transformer backbone. The backbone is then interfaces with ODE or SDE solvers [24, 59] to generate images.
@@ 5.1 Experimental Setting
We list the experimental settings below (see more details in Appendix B).
@@
For pixel space generation, we evaluate GMem on CIFAR-10 [34] for its diversity in classes and widespread use in benchmarking image generation models. We also evaluate the scalability of GMem on ImageNet $64\times 64$ [10] to assess the performance on larger resolution generation. We also conduct experiments on ImageNet $256\times 256$ [10] to evaluate the ability of GMem to model the latent space distribution, a key approach in current image generation research [28, 70, 44, 49].
Backbones.
Similar to prior image generation works [70], we utilizes SiT [44] as backbone. This architecture represent most recent advances in diffusion and flow-based transformers [70], and are widely used in image generation tasks [20, 7, 70].
@@
For a fair comparison, we compare to the SoTA image generation methods on both training efficiency and performance. Specifically, for pixel-space image generation, we consider the following three categories of baselines: First, we compare GMem with traditional generative models, including Diffusion GAN [67], Diffusion StyleGAN [65], DMD2 [69]. We also compare the SoTA diffusion models with UNets, including DDPM [24], Score SDE [61], EDM [28], DPM-Solver [43], ADM [12], EDMv2 [29], CTM [30], SiD [72], EMD [28]. Finally, we also compare to the SoTA flow-based transformer methods, including DiT [49], SiT [44], and the most recent yet contemporaneous work REPA [70].
Table 1: Sampling quality on various datasets. We report the performance of GMem on CIFAR-10 (left), ImageNet $64\times 64$ (right), and ImageNet $256\times 256$ (middle). GMem achieves comparable FID with fewer sampling steps across multiple datasets, further highlighting its advantages in accelerating training.
@@
METHOD NI ($\downarrow$) FID ($\downarrow$)
@@
Diffusion GAN [67] 51.2M 3.75
Diffusion StyleGAN [65] 50M 3.19
@@
DDPM [24] 50M 3.17
Score SDE (deep) [61] 166M 2.20
EDM [28] 200M 2.01
SiD [72] 800M 1.92
@@
SiT-XL/2 [44] 25.6M 6.68
+ REPA [70] 10M 4.52
+ GMem (ours) 10M 1.59
+ GMem (ours) 2.6M 4.08
+ GMem (ours) 22.5M 1.22
ImageNet 64$\times$64
METHOD NI ($\downarrow$) FID ($\downarrow$)
@@
SiT-XL [44]
+ GMem (ours) 22.5M 2.10
ImageNet 256$\times$256
METHOD NI ($\downarrow$) FID ($\downarrow$)
@@
MaskDiT [71] 1B 5.69
Simple diffusion [25] 1B 3.75
SD-DiT [73] 512M 5.06
DiT-XL/2 [49] 1.8B 9.60
SiT-XL/2 [44] 1.8B 8.30
+ REPA [70] 1B 5.90
+ GMem (ours) 100M 4.98
+ GMem (ours) 512M 3.56
@@
In line with prior research [44, 25], all generation quality metrics are reported as FID-50K (FID[22] scores. For CIFAR-10 dataset, we use the training set as the reference set, and for ImageNet $64\times 64$ and ImageNet $256\times 256$, we use the validation set as the reference set. Following Karras et al. [28], we also report the Number of trained Images (NI) as a measure of training efficiency. With similar FID scores, the model with less NI is considered more efficient111 $100$M NI corresponds to a total of $\sim 400$K training steps with the batch size of $256$, and baselines of different batch-sizes or training steps can be converted to NI in a similar way. . Moreover, following [42], we evaluate the efficiency of the sampling by comparing the Number of Function Evaluations (NFE) required to achieve a certain FID score. Under our setting, a lower NFE corresponds to reduced computational costs during sampling, thus making the model more efficient.
@@
We strictly adhere to the setup details outlined in [44] to ensure a fair comparison. For latent-space generation, we utilize ImageNet [10], where images are pre-processed to a resolution of $256\times 256$. We follow the ADM framework [12] for additional preprocessing strategies. Followed by Yu et al. [70], the preprocessed images then been fed into a Stable Diffusion VAE [55] and embedded into the latent space $\mathbf{z}\in\mathbb{R}^{32\times 32\times 4}$. For model configurations, we employ the B/2, L/2, and XL/2 architectures, as introduced by SiT [44], for ImageNet $256\times 256$.
We consistently use the batch size of $128$ for CIFAR-10 following [61] and $256$ for ImageNet following [29] during training. Additional experimental implementation details are provided in Appendix C.
@@
Following SiT [44], we use Euler Maruyama sampler (sde solver) and set the NFE as $50$ for CIFAR-10 and $250$ for ImageNet $64\times 64$ and ImageNet $256\times 256$ by default.
5.2 Main Results
Table 2: Ablation study and sensitive analysis. All models are based on SiT-L/2 and trained for 100M NI (400K iterations under the batch size of 256) without using classifier-free guidance on ImageNet $256\times 256$. $\downarrow$ indicates lower values are better.

### v2

@@ 4.1 External Memory Bank
Building on Kadkhodaie et al. (2023)’s insight that diffusion models achieve generalization through geometry-adaptive harmonic representations, we design a memory bank for diffusion models to: (a) supply essential semantic information for generating high-quality, realistic, and distribution-aligned images, and (b) exclude excessive details to prevent overfitting to the training data while maintaining robust generalization.
We formalize the memory bank as a matrix $\mathbf{M}\in\mathbb{R}^{n\times m}$, consisting of $n$ unit-norm memory snippets:
$\mathbf{M}=\begin{bmatrix}\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{n}\end{bmatrix}^{\top}\,,\quad|\mathbf{s}_{i}|=1\,.$ (6)
We employ a representation model $\boldsymbol{f}$ such that for any input $\mathbf{x}\sim D$, the normalized feature $\nicefrac{{\boldsymbol{f}(\mathbf{x})}}{{\|\boldsymbol{f}(\mathbf{x})\|}}$ corresponds to a vector $\mathbf{s}\in\mathbf{M}$. Through optimization, we ensure that $\mathbf{M}$ fully captures the semantic information of $D$.
Representation models for memory bank construction.
GMem supports diverse representation models $\boldsymbol{f}$, each varying in their ability to capture image information, which also directly influences generative model performance. We employ self-supervised representation models for two key reasons:
@@
    Yu et al. (2024) demonstrate that self-supervised models can expedite the training of diffusion models. It motivates to use it in GMem.
@@
    Self-supervised models capture semantic information more effectively than supervised alternatives (Bordes et al., 2022; Zimmermann et al., 2021; Sun et al., 2024).
We examine multiple models, including CLIP visual encoder (Radford et al., 2021), to construct the memory bank (see Appendix E).
Reducing network dependency on memory bank.
Directly using memory snippets to guide network generation can lead to overfitting, as the network may rely too heavily on the strong guidance provided, impairing its ability to generalize to unseen samples (Blattmann et al., 2022). To address this issue, we propose a feature-level random masking strategy. Instead of utilizing the complete memory snippets during training, we randomly mask a subset of dimensions by setting them to zero. For instance, given a memory snippet with 512 dimensions, we can randomly select 256 dimensions to mask.
We will detail the performance of various masking strategies and ratios in Section 5.4 and discuss how to select the most appropriate masking strategy.
@@ 4.2 Training Objective
Memory bank provides semantic information about the data distribution, aiding both training and inference phases in diffusion models. To integrate this, we adapt the training loss of diffusion models as follows:
@@
where $\mathbf{x}_{0}\sim D,\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),\dot{\alpha}_{t}=\frac{\mathrm{d}\alpha_{t}}{\mathrm{d}t},\dot{\sigma}_{t}=\frac{\mathrm{d}\sigma_{t}}{\mathrm{d}t}$ and $\mathbf{v}_{\boldsymbol{\theta}}$ is the velocity estimated by SiT backbone. For fair comparison, we retain the same $\alpha_{t}$ and $\sigma_{t}$ settings as in previous work (Ma et al., 2024; Yu et al., 2024).
@@ 4.3 Sampling with Memory Bank
The sampling stage of diffusion models (Ho et al., 2020; Lu & Song, 2024) transforms Gaussian noise into the true data distribution through an end-to-end framework, utilizing SDE or ODE solvers (see Section 3) to iteratively refine the noise. Similarly, GMem, equipped with a memory bank, follows this process by feeding noise and corresponding memory snippet $\mathbf{s}$ into the neural network.
However, akin to Blattmann et al. (2022); Casanova et al. (2021), the memory bank requires explicitly storage and retrieval during inference. For large-scale datasets like ImageNet, this results in prohibitive storage demands and retrieval costs. Additionally, memory snippets often exhibit significant redundancy (e.g., snippets from the same class are highly correlated), leading to unnecessary storage when all features are retained directly.
To reduce redundancy in storing memory snippets and minimize storage, we propose an efficient storage strategy.
Decomposing the memory bank into two compact matrices.
Let $\mathbf{M}\in\mathbb{R}^{N\times d}$ represent $N$ memory snippets $\mathbf{s}$, each of dimension $d$. We first compute the mean vector $\boldsymbol{\mu}\in\mathbb{R}^{d}$ and centralize $\mathbf{M}$ as $\mathbf{M}_{\text{centered}}=\mathbf{M}-\boldsymbol{\mu}$. Applying SVD (Eckart & Young, 1936) to it yields $\mathbf{M}_{\text{centered}}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$, where $\mathbf{U}\in\mathbb{R}^{N\times r}$, $\mathbf{\Sigma}\in\mathbb{R}^{r\times r}$, and $\mathbf{V}\in\mathbb{R}^{d\times r}$, with $r<\min(N,d)$ as the target latent dimension.
By distributing the singular values, we form a coefficient matrix $\mathbf{C}=\mathbf{U}\,\mathbf{\Sigma}^{1/2}\in\mathbb{R}^{N\times r}$ and a basis matrix $\mathbf{B}=\mathbf{V}\,\mathbf{\Sigma}^{1/2}\in\mathbb{R}^{d\times r}$. This results in a low-rank approximation:
$\mathbf{M}\leftarrow\mathbf{C}\,\mathbf{B}^{\top}\,,$ (8)
reducing storage complexity from $\mathcal{O}(Nd)$ to $\mathcal{O}(Nr+dr)$.
The matrix $\mathbf{B}$ acts as a compact, fixed basis encoding global structure, whereas $\mathbf{C}$ flexibly stores snippet-specific coefficients. During inference, retrieving a snippet involves looking up its coefficients in $\mathbf{C}$ and transforming them via $\mathbf{B}$. Specifically, the $i$-th memory snippet $\mathbf{s}_{i}$ can be reconstructed as:
$\mathbf{s}_{i}=\mathbf{c}_{i}\mathbf{B}^{\top}+\boldsymbol{\mu}\,,$ (9)
where $\mathbf{c}_{i}\in\mathbb{R}^{1\times r}$ is the coefficient vector corresponding to the $i$-th snippet from coefficient matrix $\mathbf{C}$.
4.4 Modular Memory Bank Enables Generalization
The core insight of GMem lies in the decoupling of memorization and generalization in a generative model, enabling generalization beyond training data. This is achieved through two approaches: (I) incorporating external knowledge by introducing novel images absent from the original dataset, and (II) manipulating internal knowledge by combining existing memory snippets into new compositions. While our memory bank provides compact storage and flexibility for new snippets integration, we note that advancing the modularity of the memory bank remains future work.
Aspect I: external knowledge manipulation.
To incorporate external knowledge outside the training dataset, we project the feature vector $\boldsymbol{f}(\mathbf{x}_{\text{new}})\in\mathbb{R}^{d}$ onto the existing coefficient matrix $\mathbf{C}$. We calculate the coefficients $\mathbf{c}_{\text{new}}\in\mathbb{R}^{r}$ of the centered feature vector by projecting it onto the basis matrix $\mathbf{B}$:
$\mathbf{c}_{\text{new}}=\nicefrac{{(\boldsymbol{f}(\mathbf{x}_{\text{new}})-\boldsymbol{\mu})\,\mathbf{B}}}{{\mathbf{S}}}\,,$ (10)
where $\mathbf{S}$ (the diagonal of $\mathbf{\Sigma}$) stores singular values [^2] These singular values are several floating-point numbers with negligible storage cost.. By appending $\mathbf{c}_{\text{new}}$ to $\mathbf{C}$ , we expanding the memory bank with negligible overhead. This process seamlessly integrates new snippets not present in the training dataset, allowing the model to utilize the network’s generalization capabilities for generating new samples without additional training.
Aspect II: internal knowledge manipulation.
We generate new memory snippets by interpolating between existing ones. Given two snippets indexed by $i$ and $j$, we construct a new coefficient vector:
$\mathbf{c}_{\text{new}}=\alpha\,\mathbf{c}_{i}+(1-\alpha)\,\mathbf{c}_{j},\quad\alpha\in[0,1]\,.$ (11)
Appending $\mathbf{c}_{\text{new}}$ to $\mathbf{C}$ yields a latent interpolation in $\mathbf{M}$ without modifying $\mathbf{B}$. This approach enables training-free style transfer and compositional generalization by exploring linear paths between coefficients of different memory snippets, effectively creating novel samples from the internal knowledge encoded in $\mathbf{C}$.
In Section 5.3, we demonstrate the effectiveness of both exact projection and interpolation for expanding the memory bank, enhancing the diversity and flexibility of generated results.
@@ 5.1 Experimental Setting
We list the setup below (see more details in Appendix A).
@@
For pixel-level generation, we test GMem on CIFAR-10 (Krizhevsky et al., 2009) due to its diverse classes and its popularity in benchmarking image generation. We then evaluate GMem on ImageNet $256\times 256$ (Deng et al., 2009) to examine how it models latent space distributions, which is a key focus in recent image generation research (Karras et al., 2022; Yu et al., 2024; Ma et al., 2024; Peebles & Xie, 2023). Finally, we assess the scalability of GMem to larger resolutions by conducting experiments on ImageNet $512\times 512$ (Deng et al., 2009).
Backbones and visual tokenizers.
Following prior image generation approaches (Yu et al., 2024), we primarily use LightningDiT (Yao & Wang, 2025) and SiT (Ma et al., 2024) as the backbone. We also evaluate the effectiveness of GMem on different visual tokenizer such as SD-VAE (Rombach et al., 2022), DC-AE (Chen et al., 2024a), and VA-VAE (Yao & Wang, 2025) in  Table 5.
These architectures represent the latest advancements in diffusion and flow-based transformers (Yu et al., 2024) and are widely adopted in image generation tasks (Han et al., 2023; Chen et al., 2024b; Yu et al., 2024).
@@
For a fair comparison, we compare to the SoTA image generation methods on both training efficiency and performance. Specifically, for pixel-space image generation, we consider the following three categories of baselines: First, we compare GMem with traditional generative models, including Diffusion GAN (Xiao et al., 2021), Diffusion StyleGAN (Wang et al., 2022), DMD2 (Yin et al., 2024). We also compare the SoTA diffusion models with UNets, including DDPM (Ho et al., 2020), Score SDE (Song et al., 2020b), EDM (Karras et al., 2022), DPM-Solver (Lu et al., 2022), ADM (Dhariwal & Nichol, 2021), EDMv2 (Karras et al., 2024), CTM (Kim et al., 2023), SiD (Zhou et al., 2024), EMD (Karras et al., 2022). Finally, we also compare to the SoTA flow-based transformer methods, including DiT (Peebles & Xie, 2023), SiT (Ma et al., 2024), and the most recent yet contemporaneous work REPA (Yu et al., 2024).
Table 1: Sampling quality on various datasets. We report the performance of GMem on CIFAR-10 (left), ImageNet $256\times 256$ (middle) and ImageNet $512\times 512$ (right). GMem achieves comparable FID with fewer sampling steps across multiple datasets, further highlighting its advantages in accelerating training. All the results reported are w/o classifier-free guidance unless otherwise specified.
@@
METHOD Epoch ($\downarrow$) FID ($\downarrow$)
@@
BigGAN (Brock, 2018) - 14.7
StyleGAN2 (Karras et al., 2020) 128 8.32
@@
DDPM (Ho et al., 2020) 100 3.17
DDIM (Song et al., 2020a) - 4.04
Score SDE (deep) (Song et al., 2020b) 333 2.20
EDM (Karras et al., 2022) 400 2.01
Diffusion Style-GAN (Wang et al., 2022) - 3.19
Diffusion GAN  (Xiao et al., 2021) 1024 3.75
@@
SiT-XL/2 (Ma et al., 2024) 512 6.68
  + REPA (Yu et al., 2024) 200 4.52
  + GMem (ours) 52 4.08
  + GMem (ours) 200 1.59
  + GMem (ours) 450 1.22
ImageNet 256$\times$256
METHOD Epoch ($\downarrow$) FID ($\downarrow$)
@@
MaskGIT (Chang et al., 2022) 300 6.18
MAGVIT-v2 (Yu et al., 2023) 270 3.65
SD-DiT (Zhu et al., 2024) 480 7.21
DiT-XL/2 (Peebles & Xie, 2023) 1400 9.62
SiT-XL/2 (Ma et al., 2024) 1400 8.30
  + REPA (Yu et al., 2024) 782 5.90
FasterDiT (Yao et al., 2024) 400 7.91
LightningDiT (Yao & Wang, 2025) 800 2.17
  + GMem (ours) 160 1.53
ImageNet 512$\times$512
METHOD Epoch ($\downarrow$) FID ($\downarrow$)
Traditional generative models
StyleGAN-XL (Sauer et al., 2022) - 2.41
BigGAN (Brock, 2018) 472 9.54
@@
MaskGIT (Chang et al., 2022) 300 7.32
MAGVIT-v2 (Yu et al., 2023) 270 3.07
DiT-XL (Peebles & Xie, 2023) 600 12.03
SiT-XL (Ma et al., 2024) (w/ cfg) 600 2.62
  + REPA (Yu et al., 2024) (w/ cfg) 200 2.08
LightningDiT-XL (Yao & Wang, 2025) - -
  + GMem (ours) 200 1.93
  + GMem (ours) 400 1.89
  + GMem (ours) (w/ cfg) 400 1.71
@@
[…]
@@
Unless otherwise specified, we adhere to the setup details outlined in (Ma et al., 2024) to ensure a fair comparison. For latent-space generation, we utilize ImageNet $256\times 256$ and $512\times 512$. We follow the ADM framework (Dhariwal & Nichol, 2021) for additional preprocessing strategies. Following Ma et al. (2024), preprocessed images are fed into SD-VAE (Rombach et al., 2022) and embedded into the latent space $\mathbf{z}\in\mathbb{R}^{32\times 32\times 4}$. Additionally, we add representation alignment loss intrfoduced from REPA (Yu et al., 2024) to help the model learn more structured representations across all experiments. For model configurations, unless otherwise specified, we empolys SiT-XL for CIFAR-10 and LightningDiT-XL for ImageNet $256\times 256$ and $512\times 512$.
We consistently use the batch size of $128$ for CIFAR-10 following (Song et al., 2020b) and $1024$ for ImageNet following (Karras et al., 2024) during training. Additional experimental implementation details are provided in Appendix B.
@@
Following SiT (Ma et al., 2024), we use SDE solver and set the NFE as $50$ for CIFAR-10 and $100$ for ImageNet $256\times 256$ and $512\times 512$ by default. More details are provided in Appendix B.
Figure 4: Demonstration of novel and compositional image generation via memory bank manipulation. Selected samples from ImageNet $256\times 256$ generated by the GMem. In the “Novel image generation” part, we show the reference image used to build a new memory snippet (left), followed by the generated samples and 5 of the nearest training images, illustrating GMem’s adaptation to external knowledge. In the “Compositional image generation” examples, two reference images (left and right) form an interpolated image (center), demonstrating GMem can manipulate internal knowledge to create new concepts.
Figure 5: Demonstration of diverse generation by GMem. Selected samples from ImageNet $256\times 256$ generated by the GMem. This figure demonstrates the diversity of images generated by GMem, which differ from the original training set in form, style, and color. This shows that GMem does not simply memorize images from the training set, but rather generates novel variations.
5.2 Training and sampling efficiency of GMem
Table 2: GMem offers 50$\times$ training speedup compared to SiT. This table compares training speedup provided by GMem to other baselines on ImageNet $256\times 256$. $\downarrow$ indicates lower is better, and all results are without classifier-free guidance.